OpenStudy (anonymous):

Which expression is a simplification of ((x^2+7x+10)/(x^2+2x-3))/(x^2+4x-5)/(x^2-x-12))? F.(x+4)/X+4 G.x+5/(x+3)(x-4) H. x+2 /x+4 I.(x+2)(x-4)/(x-1)(x-1)

OpenStudy (anonymous):

factor each polynomial invert and multiply, the rational function in the denominator

OpenStudy (anonymous):

I forgot how to do eveyrthing.... and what a lot fo those words means sadly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

summer hehe

OpenStudy (anonymous):

one big one is to take the fraction on the bottom and flip it over to multiply times the first fraction\[\frac{x ^{2}+7x+10}{x ^{2}+2x-3} \frac{x ^{2}-x-12}{x ^{2}+4x-5}\]

OpenStudy (anonymous):

now if you factor each quadratic (ax^2+bx+c), then you will be in a good position to cancel things

OpenStudy (anonymous):

To factor... take one of the quadratics and look at the last number\[x ^{2}+7x+10\] now find two numbers that when multiplied equal that number at the end... 10

OpenStudy (anonymous):

1X10, 2X5

OpenStudy (anonymous):

do i do it for both sides?

OpenStudy (anonymous):

if we add those in (both positive or both negative) we get 1+10=11 or 2+5+7 you will need to do it for each quadratic in the problem

OpenStudy (anonymous):

I find that 2X5 equals 10, which is the end number. While 2+5 is 7, which is the middle number. So my quadratic factors into (x+2)(x+5).

OpenStudy (anonymous):

so would it be? \[(x+2)(x+5)divx ^{2}+2-3 * (X+3)(x-4)divX ^{2}+4X-5\]

OpenStudy (anonymous):

now I have\[\frac{(x+2)(x+5)}{x ^{2}+2x-3} \frac{x ^{2}-x-12}{x ^{2}+4x-5}\]yeah, you got the numerator of the second one too! :)

OpenStudy (anonymous):

now get the denominators, and start cancelling out the ones in common between the top and the bottom

OpenStudy (anonymous):

to make the fractions nice in the editor use frac{numerator}{denominator}

OpenStudy (anonymous):

It looks like you're trying to make it prettier, so I thought I would give a suggestion. Have you factored the denominators?

OpenStudy (anonymous):

SORRY I WAS in the bathrookm thanks for your patience!!

OpenStudy (anonymous):

no problem... I do that too! :)

OpenStudy (anonymous):

okay i think I factored the denominators

OpenStudy (anonymous):

what did you get... you can show me the denominators, or what you get after cancelling common factors.

OpenStudy (anonymous):

I got \[(x+2)(x+5)\div(x-1)(x+3) * (x+3)(x-4)\div(x-1)(x+5)\]

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

now cancel the ones that appear both above and below... (x+5) and (x+3), the answer will be what is left behind.

OpenStudy (anonymous):

so: (x+2)(x-4)/(x-1)(x-1) is what i got

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

yay!! thanks you!!

OpenStudy (anonymous):

You're welcome. Now I have to make dinner for the family.... see ya around.

OpenStudy (anonymous):

oh! thank you so much so much responsibility but you still help em i really appreaciate it:)

OpenStudy (anonymous):

no problem... click best response button to give me a point ;)

OpenStudy (anonymous):

woof!